FOURIER TRANSFORM INFRARED EMISSION SPECTRA OF BeH/BeD AND BeH_{2}/BeD_{2}

\underline{A. SHAYESTEH}, K. TERESZCHUK, and P. F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, ON, N2L 3G1, Canada; R. COLIN, Laboratoire de Chimie Physique Moleculaire, Universite Libre de Bruxelles, C.P. 160/09, 50 av. F.D. Roosevelt, 1050 Brussels, Belgium.

We have recorded the first vibration-rotation spectra of gaseous BeH/BeD and BeH_{2}/BeD_{2} in emission using a furnacedischarge source. The molecules were generated at 1500^{o}C and 333 mA discharge current with beryllium and a mixture of helium and hydrogen or deuterium gases. The recorded spectra contained several emission bands, as well as the absorption of atmospheric H_{2}O and CO_{2}. The highest signal-to-noise ratio for BeH lines was about 350. The vibrational bands v=1-0 to v=4-3 of BeH/BeD were observed in the spectra. The bands were rotationally analyzed, and the spectroscopic constants for v=0 to 4 levels of BeH and BeD, as well as the Dunham constants and the Born-Oppenheimer breakdown constants, were obtained. The antisymmetric stretching mode (ν_{3}) of BeH_{2}/BeD_{2} and several hot bands involving ν_{1}, ν_{2} and ν_{3} were found and rotationally analyzed to yield the spectroscopic constants. For BeH and BeD, we determined the equilibrium rotational constants (B_{e}) to be 10.31959(3) cm^{−1} and 5.68829(2) cm^{−1}, respectively, while the equilibrium vibrational constants (ω_{e}) are 2061.416(3) cm^{−1} and 1529.956(3) cm^{−1}. We also determined B_{e} for BeH_{2} to be 4.75366(2) cm^{−1}. Using the B_{e} values of BeH and BeH_{2}, the equilibrium Be-H bond distances (R_{e}) were calculated for BeH and BeH_{2} to be 1.342436(2) Å and 1.326407(3) Å, respectively.